Write a system of two equations in two unknowns. Solve the system by using the substitution method. « The length of a rectangular field is 12 feet longer than the width. If the perimeter is 144 feet, then what are the length and width
Answer by rfer(16322)

y=x+12
2x+2y=144
--------------
2x+2(x+12=144
2x+2x+24=144
4x=120
x=30
---------------
y=42
--------------
proof
2(30)+2(42)=144
60+84=144